OpenStudy (anonymous):

Choose the correct simplification of the expression -5x2(4x - 6x2 - 3). (4 points) A) 30x4 - 20x3 + 15x2 B) -11x4 - x3 - 8x2 C) -30x4 + 20x3 - 15x2 D) 30x4 + 20x3 + 15x2

OpenStudy (anonymous):

@mathmath333 @mrdoldum Help?

OpenStudy (anonymous):

multiply each term inside the parenthesis with \[-5x^2\] and add them

OpenStudy (anonymous):

\[a(b+c)=ab+ac\]

OpenStudy (anonymous):

So it would be A? :/

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ -5x^2(4x - 6x^2 - 3). \hspace{.33em}\\~\\ =\color{red }{-5x^2\times 4x} - (-5x^2)6x^2 -(-5x^2) 3 \hspace{.33em}\\~\\ }\end{align}\) \(\large \begin{align} \color{black}{ -5x^2\times 4x \hspace{.33em}\\~\\ =-5\times 4\times x^2\times x^1 \hspace{.33em}\\~\\ =-20\times x^{2+1} \hspace{.33em}\\~\\ =-20\times x^{3} \hspace{.33em}\\~\\ }\end{align}\) like this simplify the other two terms

OpenStudy (anonymous):

That's something you must find

OpenStudy (anonymous):

okay. Thank you both(:

OpenStudy (beccasmith):

Sorry i just really need help

OpenStudy (beccasmith):

I will delete the comment

OpenStudy (anonymous):

then go tag someone in your department

OpenStudy (anonymous):

@beccasmith Out of principle alone I wouldn't help a person doing what you are doing here. You need to make your own thread, not try to hijack some one else's. It is rude. How would you like it if some one did that to your thread?

OpenStudy (beccasmith):

Honestly I would look at theirs and not really care.